3x^2+12x+5=p(x+q)^2+r
find values for p,q and r
I appreciate any form of help you can give me.

Answer by uma(370)

3x^2+12x+5=p(x+q)^2+r
Consider the left side:
3x^2 + 12x + 5
= 3(x^2 + 4x) + 5
= 3(x^2 + 4x + 4 - 4) + 5 [ add and subtract to make it a perfect square]
=3(x^2 + 4x + 4) - 12 + 5
= 3(x+2)^2 - 7
= p(x+q)^2 + r
Equating, we get, p = 3, q = 2, r = -7
good luck!!!